An automobile is moving at speed of 2.00 m/s along a circular road of radius 20 m.

A) Find the time it takes to make one full circle when the car is still moving at constant speed. (1P)

B) Find the car radial acceleration during this time. (1P) Then its speed starts increasing at a rate of 0.600 m/s2 while it stays on the same road.

C)How long would it take for the car to make one full turn from the moment it started its tangential acceleration?
Answer #1

Initial speed U=2.00m/s radius. R = 20m Time taken to complete dit podt are are full circle - 29TR olt popp - le t 2T + 20 IT0.300 +² + 2t - 40T1 = 0 0 300t² & 2t - 125.66 = 0 t = 17.45)